OpenStudy (anonymous):

is this right 14% of 90 90/14=6 6*100 your answer is 600

OpenStudy (kirbykirby):

Is the question simply asking to find 14% of 90? Because 600 isn't

OpenStudy (anonymous):

yeah I can not remember how to do this

OpenStudy (kirbykirby):

90*14... then divide by 100

OpenStudy (anonymous):

do 90 *.14

OpenStudy (anonymous):

Percent means per 100, so 100% of a value is the value itself. One percent is the value times 1/100, so 14% is the value times 14/100.

OpenStudy (kirbykirby):

\[\frac{14}{100}*90=\frac{14*90}{100}\]

OpenStudy (anonymous):

so it will be 12.6 for the answer

OpenStudy (anonymous):

14 % of 90 0.14 x 90 = ...... 0.14/100 x 90 = .....

OpenStudy (anonymous):

1260

OpenStudy (anonymous):

nope, try again.

OpenStudy (anonymous):

12.6

OpenStudy (kirbykirby):

not 0.14/100... it would be 14/100 @JayDS :)

OpenStudy (anonymous):

u may have made an error typing in the values.

OpenStudy (anonymous):

no I mean 90*14 i2 1260

OpenStudy (anonymous):

woops my bad, mistyped it.

OpenStudy (anonymous):

14 % of 90 0.14 x 90 = ...... 14/100 x 90 = .....

OpenStudy (kirbykirby):

yes that is correct, now divide by 100! (Just move the decimal two places towards the left).